# BSRBF-KAN: A combination of B-splines and Radial Basis Functions in Kolmogorov-Arnold Networks

### v1

@@ Abstract
In this paper, we introduce BSRBF-KAN, a Kolmogorov Arnold Network (KAN) that combines B-splines and radial basis functions (RBFs) to fit input vectors in data training. We perform experiments with BSRBF-KAN, MLP, and other popular KANs, including EfficientKAN, FastKAN, FasterKAN, and GottliebKAN over the MNIST dataset. BSRBF-KAN shows stability in 5 training times with a competitive average accuracy of 97.55% and obtains convergence better than other networks. We expect BSRBF-KAN can open many combinations of mathematical functions to design KANs. Our repo is publicly available at: https://github.com/hoangthangta/BSRBF-KAN.
@@ 1 Introduction
With that inspiration, many scientists have flocked to develop different types of KANs based on popular polynomial functions. While these works focus on univariate functions to set up KANs, none have explored their combination. Therefore, we aim to combine B-splines [2] and Radial Basis Functions [3] to build a combined KAN named BSRBF-KAN. The reason why we use them is their successful implementation in two popular KANs, EfficientKAN [^1] https://github.com/Blealtan/efficient-kan and FastKAN [4]. Additionally, they are continuous basis functions used for function approximation, providing features such as smoothness, adaptability, and local control. With the experiments on MNIST, we show the stability and faster convergence of BSRBF-KAN compared to other networks.
@@ 2 Related Works
KANs are considered a fresh breeze that blew into the scientific community in May 2024 to design different KANs. Many works rely on popular mathematical functions, especially ones that can deal with curves such as B-Splines [2] (Original KAN [1], EfficientKAN [^2] https://github.com/Blealtan/efficient-kan), Gaussian Radial Basis Functions or GRBFs (FastKAN [4]), Reflection SWitch Activation Function or RSWAF (FasterKAN [15]), Chebyshev polynomials (TorchKAN [16], Chebyshev KAN [17]), Legendre polynomials (TorchKAN [16]), Fourier transform (FourierKAN [^3] https://github.com/GistNoesis/FourierKAN/, FourierKAN-GCF [18]) and other polynomial functions [19]. In this paper, we prefer to work on combining basic mathematical functions to design KANs instead of single functions to explore the effects of combined functions on KANs.
@@ 3.4 BSRBF-KAN
Inspired by successful implementations of previous KAN models, we developed BSRBF-KAN by combining B-splines and RBFs, which are widely recognized in computational mathematics for interpolation and approximation tasks. RBFs here we choose are Gaussian ones because they are popular. Both B-splines and RBFs are smooth functions, ensuring continuity and derivatives up to a specified order, making them suitable for choosing proper $\Phi_{q}$ and $\phi_{q,p}$ in KANs.
@@ 4 Experiments
KAN models were trained on the MNIST dataset [22] with a structure of (784, 64, 10), comprising 784 input neurons, 64 hidden neurons, and 10 output neurons corresponding to the 10 output classes (0-9). For GottliebKAN, we used a network structure (784, 64, 64, 10) to make it deeper. Other hyperparameters are the same in all models, including batch_size = 64, learning_rate = 1e-3, weight_decay=1e-4, optimize = AdamW, and loss=CrossEntropy. We conducted 5 training times for each model to assess their overall performance more reliably. By taking the best metric value from each time, we aimed to mitigate the effects of variability in training and ensure that the evaluation reflects the models’ peak capabilities.
Table 1: The best metric values in 5 training times.
Model Train. Acc. Val. Acc. F1 Time (seconds) #Params
BSRBF-KAN 100.0 97.63 97.6 222 459040
@@
Table 2: The average metric values in 5 training times.
Model Train. Acc. Val. Acc. F1 Time (seconds)
BSRBF-KAN 100.00 ± 0.00 97.55 ± 0.03 97.51 ± 0.03 231
@@
Table 1 shows the best results over 5 training times. Due to its deeper network structure, GottliebKAN outperformed the other models with a validation accuracy of 97.78% and an F1 score of 97.74%. MLP and BSRBF-KAN achieved the second-best and third-best performance, though it took a long training time. Conversely, FasterKAN had the shortest training time, with its performance only surpassing that of FastKAN, which was the lowest. BSRBF-KAN is the only model obtaining a training accuracy of 100% after 15 epochs. If counting the results based on average values, MLP is the best model, as displayed in Table 2, followed competitively by BSRBF-KAN. Furthermore, both models obtain stability over 5 training times compared to the others. Despite achieving the highest accuracy values, GottliebKAN is unstable.
Figure 1: The logarithm values of training losses (left) and validation losses (right) in 15 epochs.
Figure 1 demonstrates training and validation losses of models represented by logarithm values in a certain training time. The training loss values of BSRBF-KAN drop sharply continuously, which makes it the model with the smallest loss after the training process. A similar phenomenon also happens with its validation losses, but only in 5 first epochs. The model that takes the least loss is GottliebKAN.
Through our experiments, we observed that combined models like BSRBF-KAN enhance performance and facilitate model convergence compared to single KANs, although they require more training time. This trade-off is an important consideration when selecting suitable models for different types of problems. Besides, the key benefits of using KANs, in general, are not clear when compared to MLP, at least in this paper.
5 Ablation Study
We conducted several ablation experiments to evaluate the impact of various components within BSRBF-KAN on overall model performance. Table 3 presents the metric values for different combinations of BSRBF-KAN components. Initially, we trained the model with all components included, followed by training variants with some components removed. The results show that the base output and layer normalization are key for the model’s performance, as their absence leads to the worst results.
While B-splines and RBFs can enhance the performance, removing one has minimal impact on the model. Interestingly, when both are absent, the model becomes an MLP, which shows slightly better validation accuracy and F1 scores than the complete BSRBF-KAN model. However, it has less model convergence ability compared to BSRBF-KAN.
Table 3: BSRBF-KAN by different components.
Components Train. Acc. Val. Acc. F1
Full 100.0 97.53 97.49
No RBF - 0.01 - 0.14 - 0.14
No BS - 0.02 - 0.21 - 0.21
No BS + No RBF = MLP - 0.55 + 0.03 + 0.06
@@ 6 Conclusion
We introduced BSRBF-KAN, a KAN that integrates B-splines and RBFs to enhance stability and data training convergence. We did not intend to make BSRBF-KAN better than MLP and other KANs by optimizing hyperparameters; instead, we built it by default to understand aspects of KANs combination in training models. In the experiments, we trained BSRBF-KAN, MLP, and other KANs on the MNIST dataset. The results revealed that while GottliebKAN achieved the highest accuracy, it lacked stability across 5 training times. MLP and BSRBF-KAN demonstrated stability and competitive performance, with BSRBF-KAN showing better model convergence than MLP. In the future, we will continue to exploit the combination of KANs and their components and design KAN’s architecture to improve the model performance in broader problems and datasets.

### v2

@@ Abstract
In this paper, we introduce BSRBF-KAN, a Kolmogorov Arnold Network (KAN) that combines B-splines and radial basis functions (RBFs) to fit input vectors in data training. We perform experiments with BSRBF-KAN, MLP, and other popular KANs, including EfficientKAN, FastKAN, FasterKAN, and GottliebKAN over the MNIST and Fashion-MNIST datasets. BSRBF-KAN shows stability in 5 training sessions with a competitive average accuracy of 97.55% on MNIST and 89.33% on Fashion-MNIST and obtains convergence better than other networks. We expect BSRBF-KAN to open many combinations of mathematical functions to design KANs. Our repo is publicly available at: https://github.com/hoangthangta/BSRBF-KAN.
@@ 1 Introduction
With that inspiration, many scientists have flocked to develop different types of KANs based on popular polynomial and basis functions. While these works focus on univariate functions to set up KANs, none have explored their combination. Therefore, we aim to combine B-splines [2] and Radial Basis Functions [3] to build a combined KAN named BSRBF-KAN. The reason why we use them is their successful implementation in two popular KANs, EfficientKAN [^1] https://github.com/Blealtan/efficient-kan and FastKAN [4]. Additionally, they are continuous basis functions used for function approximation, providing features such as smoothness, adaptability, and local control. With the experiments on MNIST [5] and Fashion-MNIST [6], we show the stability and faster convergence of BSRBF-KAN compared to other networks.
@@ 2 Related Works
KANs are a fresh breeze that has blown into the scientific community since May 2024. Many KAN design works rely on popular mathematical functions, especially ones that can deal with curves such as B-Splines [2] (Original KAN [1], EfficientKAN [^2] https://github.com/Blealtan/efficient-kan), Gaussian Radial Basis Functions or GRBFs (FastKAN [4]), Reflection SWitch Activation Function or RSWAF (FasterKAN [17]), Chebyshev polynomials (TorchKAN [18], Chebyshev KAN [19]), Legendre polynomials (TorchKAN [18]), Fourier transform (FourierKAN [^3] https://github.com/GistNoesis/FourierKAN/, FourierKAN-GCF [20]) and other polynomial functions [21]. In this paper, we prefer to work on combining basic mathematical functions to design KANs instead of single functions to explore the effects of combined functions on KANs.
@@ 3.4 BSRBF-KAN
Inspired by successful implementations of previous KAN models, we developed BSRBF-KAN by combining B-splines and RBFs, which are widely recognized in computational mathematics for interpolation and approximation tasks. The RBFs we chose here are Gaussian ones because they are popular. Both B-splines and RBFs are smooth functions, ensuring continuity and derivatives up to a specified order, making them suitable for choosing proper $\Phi_{q}$ and $\phi_{q,p}$ in KANs.
@@ 4 Experiments
We conducted 5 training sessions for each model on MNIST [5] and Fashion-MNIST [6] to assess their overall performance more reliably. By taking the best metric value from all sessions, we aimed to mitigate the effects of variability in training and ensure that the evaluation reflects the models’ peak capabilities. With the intention of designing simple networks, we only use activation functions, linear transformation, and layer normalization.
All models have a network structure of (784, 64, 10), comprising 784 input neurons, 64 hidden neurons, and 10 output neurons corresponding to the 10 output classes (0-9). For GottliebKAN, we used a structure (784, 64, 64, 10) to make it the same as the original design [^5] https://github.com/seydi1370/Basis_Functions. We trained the models with 15 epochs on MNIST and 25 epochs on Fashion-MNIST. Other hyperparameters are the same in all models, including batch_size = 64, learning_rate = 1e-3, weight_decay=1e-4, optimize = AdamW, and loss=CrossEntropy.
Figure 1 demonstrates training losses of models represented by logarithm values in a certain training time. On both MNIST and Fashion-MNIST, the training loss values of BSRBF-KAN drop sharply continuously, which makes it the model with the smallest loss after the training process. In contrast, GottliebKAN lost the least loss, which affected its performance in the next experiments.
Figure 1: The logarithm values of training losses (left) and validation losses (right) in 15 epochs in a training session.
Table 1: The best metric values in 5 training sessions on MNIST and Fashion-MNIST.
Dataset Model Train. Acc. Val. Acc. F1 Time (seconds) #Params
MNIST BSRBF-KAN 100.0 97.63 97.6 222 459040
@@
Fashion-MNIST BSRBF-KAN 99.3 89.59 89.54 219 459040
FastKAN 98.27 89.62 89.6 160 459114
FasterKAN 94.4 89.39 89.3 157 408224
EfficientKAN 94.83 89.11 89.04 182 508160
GottliebKAN 93.79 87.69 87.61 241 219927
MLP 93.58 88.51 88.44 147 52512
Average of MNIST + Fashion-MNIST BSRBF-KAN 99.65 93.61 93.57 220.5 459040
FastKAN 99.11 93.50 93.47 131 459114
FasterKAN 96.46 93.39 93.33 125 408224
EfficientKAN 97.09 93.33 93.27 152 508160
GottliebKAN 96.73 92.74 92.68 255 219927
MLP 96.50 93.10 93.05 210 52512
@@
Table 2: The average metric values in 5 training sessions on MNIST and Fashion-MNIST.
Dataset Model Train. Acc. Val. Acc. F1 Time (seconds)
MNIST BSRBF-KAN 100.00 ± 0.00 97.55 ± 0.03 97.51 ± 0.03 231
@@
Fashion-MNIST BSRBF-KAN 99.19 ± 0.03 89.33 ± 0.07 89.29 ± 0.07 211
FastKAN 98.19 ± 0.04 89.42 ± 0.07 89.38 ± 0.07 162
FasterKAN 94.40 ± 0.01 89.26 ± 0.06 89.17 ± 0.07 154
EfficientKAN 94.76 ± 0.06 88.92 ± 0.08 88.85 ± 0.09 183
GottliebKAN 90.66 ± 1.08 87.16 ± 0.24 87.07 ± 0.25 238
MLP 93.56 ± 0.05 88.39 ± 0.06 88.36 ± 0.05 148
Average of MNIST + Fashion-MNIST BSRBF-KAN 99.60 93.44 93.40 221
FastKAN 99.07 93.34 93.30 131.5
FasterKAN 96.44 93.27 93.21 123.5
EfficientKAN 97.07 93.15 93.09 151.5
GottliebKAN 94.55 92.18 92.11 229.5
MLP 96.50 93.01 92.98 164.5
@@
4.1 MNIST results analysis
As shown in Table 1, GottliebKAN is better than the other models, with a validation accuracy of 97.78% and an F1 score of 97.74%. We assume this because of its deeper network structure. MLP and BSRBF-KAN achieved second-best and third-best performances but took a long time to train. Conversely, FasterKAN had the shortest training time, but its performance slightly surpassed FastKAN. BSRBF-KAN is the only model with a training accuracy of 100% after 15 epochs. If counting the results based on average values, MLP is the best model, as displayed in Table 2, followed competitively by BSRBF-KAN. Despite achieving the highest accuracy values, GottliebKAN is unstable.
4.2 Fashion-MNIST results analysis
The ease of achieving high accuracy of models trained on MNIST may not accurately reflect their actual performance. Thus, we experimented with the model training on Fashion-MNIST. As shown in Table 1 and Table 2, FastKAN obtained the highest validation accuracy. At the same time, BSRBF-KAN followed competitively with and still kept the best convergence ability. MLP is the second-worst model, only performing better than GottliebKAN, but it has the fastest training time and the fewest parameters. However, when considering training time, we see that FasterKAN’s training time is only slightly longer than MLP’s, while achieving nearly 0.9% higher validation accuracy, according to the results in Table 1.
4.3 MNIST + Fashion-MNIST results analysis
When calculating the average values on MNIST and Fashion-MNIST, we clearly see that BSRBF-KAN achieved the best performance and also the best convergence, according to Table 1 and Table 2. Furthermore, all KANs except for GottliebKAN surpassed MLP. This indicates the positive signal of using KANs to obtain better accuracy values in model training, at least in these experiments.
4.4 Ablation Study
Several ablation experiments were performed to help us understand the impact of various components within BSRBF-KAN on overall model performance. Table 3 presents the metric values for different combinations of BSRBF-KAN components. Initially, we trained the model with all components included, followed by training variants with some components removed. The results show that the base output and layer normalization are key for the model’s performance, as their both absence leads to the worst results.
While B-splines and RBFs can enhance performance, removing RBFs have less impact on the model than B-splines. Interestingly, when both are absent, the model becomes an MLP, which exhibits slightly better validation accuracy and F1 scores than the complete BSRBF-KAN model on MNIST, but on Fashion-MNIST, this is pretty bad. Thus, we believe that retaining B-splines and RBFs remains beneficial for the model’s performance.
Table 3: The performance of BSRBF-KAN by different components on MNIST and Fashion-MNIST in a training setup.
Dataset Components Train. Acc. Val. Acc. F1
MNIST Full 100.0 97.53 97.49
No BS - 0.01 - 0.21 - 0.21
No RBF - 0.003 - 0.14 - 0.14
No BS + No RBF = MLP - 0.55 + 0.05 + 0.06
@@ 5 Conclusion
We introduced BSRBF-KAN, a novel KAN that integrates B-splines and Radial Basis Functions (RBFs) to fit input vectors in the training process. We did not intend to make BSRBF-KAN better than MLP and other KANs by optimizing hyperparameters; instead, we built it by default to understand aspects of KANs combination in training models. In the experiments, we trained BSRBF-KAN, MLP, and other KANs on the MNIST and Fashion-MNIST datasets. BSRBF-KAN demonstrated competitive performance and rapid convergence compared to other models. Furthermore, it is the best model if counting on the average values of both datasets. However, its high convergence rate can lead to overfitting, which can be mitigated by using dropout.
All models were generally stable, except for GottliebKAN. Although GottliebKAN achieved the highest validation accuracy on MNIST, it performed the worst on Fashion-MNIST. Similarly, while MLP achieved the highest average validation accuracy on MNIST, it underperformed compared to other KANs on Fashion-MNIST. Therefore, it is crucial to experiment with our approach on other datasets in the future to gain a more comprehensive understanding. We will also continue to explore the combination of KANs and their components when designing new KAN architectures to enhance model performance across a wider range of problems.